The average monthly salary of 86 employees has been calculated as ¢1516. Employee No. 87 earned ¢1621. What is the mean salary of the entire 87 employee?


1
Expert's answer
2020-07-05T17:29:29-0400

Solution :


Mean of n data = sum of n data / n


given mean of 86 employees = 1516

"\frac{(x1+x2+.....+x86)}{86} = 1516"


(x1+x2+.....+x86)= 1516*86

(x1+x2+.....+x86)=130376 -----------------------(i)

Mean of 87 values ="\frac{ x1+x2+.....+x86+x87}{87} = \frac{130376+ 1621}{87}"

="\frac{131997}{87}"

= 1517.21


answer: mean salary of entire 87 employees= 1517.21